Al and Cam are two cleaners. Al cleans 50 square metres of an office in 20 minutes. Cam cleans 60 square metres of the same office in 30 minutes. How long will it take Al and Cam to clean 270 square metres of this office together?
My answer is 120 minutes,but it is wrong,the correct answer is 60 minutes. Can someone show me the way how to get 60 minutes please. Thanks!

Answer by Alan3354(69443)

Al and Cam are two cleaners. Al cleans 50 square metres of an office in 20 minutes. Cam cleans 60 square metres of the same office in 30 minutes. How long will it take Al and Cam to clean 270 square metres of this office together?
--------
Al does 50/20 = 2.5 sq m/minute
Cam does 60/30 = 2 sq m/minute
------
Together they do 4.5 sq m/minute
==============
270/4.5 = 60 minutes